OpenStudy (anonymous):

Find the equation of a line parallel to the given line: y = -3/4x -1.

OpenStudy (zepp):

Parallel means same slope, in this case, the slope of the original equation is -3/4 Your equation of a line parallel to the give line would be y=-3/4(+/) Constant.

OpenStudy (zepp):

It could be y=-3/4x-9 or y=-3/4+2 etc.etc.